# Charged-Current Elastic Scattering at the Electron-Ion Collider

### v3

@@ Abstract
We discuss the measurement of the charged-current elastic scattering process $e^{-}p\rightarrow\nu_{e}n$ at the Electron-Ion Collider (EIC). This process provides sensitivity to the poorly constrained axial form factor of the nucleon, which encodes the spatial distribution of weak charge. Collisions of electrons with polarized protons enable measuring the axial form factor via the $e^{-\!}\,\vec{p}\to\nu_{e}\,n$ target-spin asymmetry for the first time. We conclude that a measurement of charged-current elastic scattering at the EIC will, perhaps unsurprisingly, prove very challenging. However, with dedicated instrumentation at a second EIC detector, the measurement may be possible.
@@ I Introduction
Understanding the sizes of nucleons and nuclei has been a primary goal of nuclear physics for over a century. The momentum-transfer dependence of elastic scattering processes provides a straightforward path to define the spatial extent of the target, and different elastic processes can provide complementary information about the distributions of quantum numbers, such as electric charge, mass, and even spin Hofstadter1956; PolyakovSchweitzer2018; Miller:2019DefiningRp; Chen:2024oxx; KaiserWeise:2024Sizes; Goharipour:2025yxm.
@@
The elastic process we will focus on in this paper is the charged-current elastic (CCE) scattering process, $e^{-}p\rightarrow\nu_{e}n$ (Fig. 1), whereby the structure of the proton is observed through the lens of a $W^{-}$. The fact that in this process the $W^{-}$ nominally couples only to left-handed up-type quarks in the proton means that it serves as a unique filter for studying flavor- and spin-dependent contributions to the nucleon structure. The form factor associated with this process is the axial form factor [^1] Also called the nucleon axial-vector form factor; some literature also uses “nucleon transition axial form factor”, since technically it describes a transition from $n\rightarrow p$ or vice versa., $F_{A}(t=-Q^{2})$ [^2] In elastic scattering, the momentum transfer to the proton and the lepton are equivalent. Therefore, $|t|$ and $Q^{2}$ can be used interchangeably, and we use only $|t|$ for convenience in the following text. LlewellynSmith1972. This form factor parametrizes the spatial distribution of axial currents in the nucleon. The so-called nucleon “axial radius”, $r_{A}$, is defined as
$r_{A}^{2}\equiv-\frac{6}{F_{A}(0)}\frac{\mathrm{d}F_{A}}{\mathrm{d}|t|}\big|_{|t|=0},$ (1)
@@
where $F_{A}^{\prime}(0)$ and $F_{A}(0)$ refer to the slope and the value of the form factor at zero momentum transfer, respectively Chen:2024oxx. The authors of Refs. Chen:2024oxx; Chen:2024ksq highlight $r_{A}$ as the primary component of a proton “spin radius”, defined as $\langle r_{\mathrm{spin}}^{2}\rangle=r_{A}^{2}+\frac{1}{4M^{2}}(1+\frac{2F_{p}(0)}{F_{A}(0)})$, where $F_{p}(0)$ is the induced pseudoscalar form factor at zero momentum transfer. $r_{A}$ provides novel information about how spin is arranged in the nucleon at a global scale, and measurement of $r_{A}$ is an attractive goal for future experiments.
@@
The value of $F_{A}(0)$ is known to high precision from $\beta$-decay processes. The measured value is $F_{A}(0)=-1.2723\pm 0.0023$, which sets the normalization of the form factor. The axial form factor is also a common target of lattice calculations Alexandrou:2017hac; Alexandrou:2020okk; Capitani:2017qpc; Gupta:2017dwj; RQCD:2019jai; Tsuji:2025quu. For the purpose of extracting the axial radius of the nucleon, the necessary quantity to measure is the slope of the form factor as $t\rightarrow 0$, denoted here as $F^{\prime}_{A}(t\rightarrow 0)$.
@@
Figure 1: Diagram of the CCE process. In the elastic process, unlike charged-current deep inelastic scattering, the $W^{-}$ does not resolve the partonic structure of the nucleon and instead is sensitive to the axial form factor.
@@
The landscape of existing experimental measurements of $F_{A}$ includes quasi-elastic neutrino scattering on nuclei, pion electroproduction on protons, and a measurement of $\bar{\nu_{\mu}}p\rightarrow\mu^{+}n$ in the organic scintillator of the MINERvA experiment Choi:1993PRL_AxialPionElectroprod; Liesenfeld:1999PLB_AxialFromPeepiN; T2K:2020CC0piOC; AguilarArevalo:2010MiniBooNECCQE; MINERvA:2023avz. Both the neutrino data with nuclear targets and the pion electroproduction data suffer from large epistemic uncertainties and model dependence NuSTEC:2017hzk; Haberzettl:2000sm. The MINERvA data therefore provide the cleanest existing measurement of the axial form factor, and the value of $r_{A}$ extracted from that data is $0.73\pm 0.17$ fm, which is consistent with the charge radius of the proton within the uncertainties.
@@
The goal of this manuscript is to investigate the feasibility of measuring the CCE process at the EIC. The highly polarized proton beam at the EIC opens the door for a totally novel measurement of $F_{A}$ via the target-spin asymmetries of the CCE process. A polarized proton target is infeasible in neutrino scattering experiments due to the large target mass required to acquire adequate statistics for a measurement. Refs. Tomalak:2020zlv; Graczyk:2019xwg calculate the target-spin asymmetries for the crossed process $\nu n\rightarrow l^{-}p$ for both longitudinal and transverse target polarizations. These target-spin asymmetries are calculated to be very large, the longitudinal asymmetry $A_{UL}$ reaches nearly 100% at low $Q^{2}$ and the transverse asymmetry $A_{UT}$ is greater than 50% for $Q^{2}>0.5\mathrm{~GeV}^{2}$. Ref. Tomalak:2020zlv further demonstrates the sensitivity of the target-spin asymmetry to the axial form factor. A combined measurement of the target-spin asymmetry and the $|t|$-dependence of the cross section of the CCE process in $ep$ scattering therefore provides a unique way to access $F_{A}$ that is complementary to neutrino scattering.
@@ II Measurement of Charged-Current Elastic Scattering at the EIC
Figure 2: Compilation of data for the reaction $\nu_{\mu}n\rightarrow\mu^{-}p$ reproduced from Ref. Formaggio:2012cpf. By crossing symmetry, this cross section should be identical to $e^{-}p\rightarrow\nu_{e}n$, neglecting lepton-mass effects which vanish at higher energies.
@@
Figure 4: Expected event yield in 100 fb-1 of integrated luminosity for the EIC in the 10x275 GeV configuration. The bin width in $|t|$ is 0.25 GeV2.
@@ V Projections
Figure 7: Left: Projected yields as a function of $|t|$ for 500 fb-1. Center: Background-subtracted CCE cross section. Right: Projected target-spin asymmetry including uncertainty on the background asymmetry $A_{bg}$. See text for more details.
@@
Assuming a veto system capable of reducing the number of background events to 5000 per 100 fb-1, we make projections for the measurement of the CCE cross section and longitudinal target-spin asymmetry in Fig. 7. The shape of the background is assumed to have the same functional form as the signal shown in Fig. 4, but with a different dipole mass of 0.8 GeV to avoid bias from having exactly the same background and signal distribution shapes. An integrated luminosity of 500 fb-1, corresponding to 5 years of running at the design luminosity, is used for these projections.
@@
Projections for the measurement of the longitudinal target-spin asymmetry $A_{UL}$ are shown in the right panel of Fig. 7. A 70% polarization on the proton beam is assumed; the impact of the nominal 1% uncertainty on the value of the polarization Huang:2025gqx; Rathmann:2025jgp is negligible compared to the statistical uncertainty. The longitudinal target-spin asymmetry of the background process $e\vec{p}\rightarrow n+X$ is currently unknown, but its magnitude will impact how well the CCE asymmetry can be extracted. To account for a lack of knowledge in the asymmetry of the background, we assume the background has an asymmetry of $0.05\pm 0.025$ and propagate this uncertainty into the final uncertainty on the measured CCE asymmetry. The resulting uncertainty is shown in the right panel of Fig. 7. This source of uncertainty can be reduced significantly using even low-luminosity data from the early runs of the EIC to measure the $e\vec{p}\rightarrow n+X$ asymmetry. The datapoints are scattered around a nominal asymmetry of $A_{UL}=1$ according to their uncertainties to mimic a real measurement. The results on $A_{UT}$ would be the similar to those shown in Fig. 7, barring a very large $A_{UT}$ for the background process $e\vec{p}\rightarrow n+X$.
The results of Ref. Tomalak:2020zlv show that a variation of $F_{A}=(1\pm 0.2)\times F_{A}^{\mathrm{Exp.}}$, where $F_{A}^{\mathrm{Exp.}}$ is the nominal experimental value of $F_{A}$, produces a roughly 10% change in $A_{UL}$ for $|t|>0.5\mathrm{~GeV^{2}}$. Below $|t|=0.5\mathrm{~GeV^{2}}$ the asymmetry depends more weakly on $F_{A}$. Therefore, it seems unlikely that this method will produce a competitive measurement of $F_{A}$, however it can be used to cross-check existing results and theoretical understanding with a novel technique [^3] One is reminded of the proton elastic form factor puzzle JeffersonLabHallA:1999epl, where the inclusion of the polarization transfer technique revealed a large discrepancy with the standard Rosenbluth method..
@@ VII Conclusion
The charged-current elastic scattering process provides a novel means to study the spin structure of the nucleon and potentially some presently unexplored corners of BSM physics. With its high luminosity and polarized beams of electrons and protons, the EIC is, in principle, an excellent tool for studying this reaction. The cross section for CCE scattering at the EIC is expected to be around 3.5 fb assuming an equal mix of left- and right-handed electrons in the beam. Target-spin asymmetries are predicted to be on the order of 100%, with their magnitude being sensitive to the axial form factor of the nucleon. However, the presence of a dominating background from inclusive neutron photoproduction makes it highly unlikely this process can be measured at the EIC without dedicated instrumentation. If one is optimistic, a dedicated hermetic veto detector may be able to deliver the necessary background suppression, depending on the final implementation of the far-forward region and the level of machine backgrounds. Such a veto system could be incorporated into the far-forward region of a second EIC detector. Even with such a veto system, it is unlikely that the EIC will deliver a competitive measurement of the axial form-factor, however the potential exists to measure target-spin asymmetries of this process for the first time, cross-checking the expectations from theory.

### v4

@@ Abstract
We discuss the measurement of the charged-current elastic scattering process $e^{-}p\rightarrow\nu_{e}n$ at the Electron-Ion Collider (EIC). This process provides sensitivity to the poorly constrained axial radius of the nucleon. Collisions of electrons with polarized protons enable measuring the $e^{-\!}\,\vec{p}\to\nu_{e}\,n$ target-spin asymmetry for the first time. We conclude that a measurement of charged-current elastic scattering at the EIC will, perhaps unsurprisingly, prove very challenging. However, with dedicated instrumentation at a second EIC detector, the measurement may be possible.
@@ I Introduction
Elastic scattering on nucleons and nuclei has been one of the primary tools of nuclear and hadronic physics for over a century. The momentum-transfer dependence of elastic scattering processes provides a path to characterize the structure of the target, and different elastic processes can provide complementary information about the distributions of quantum numbers, such as electric charge, mass, and even spin Hofstadter1956; PolyakovSchweitzer2018; Miller:2019DefiningRp; Chen:2024oxx; KaiserWeise:2024Sizes; Goharipour:2025yxm.
@@
The elastic process we will focus on in this paper is the charged-current elastic (CCE) scattering process, $e^{-}p\rightarrow\nu_{e}n$ (Fig. 1), whereby the structure of the proton is observed through the lens of a $W^{-}$. The fact that in this process the $W^{-}$ nominally couples only to left-handed up-type quarks in the proton means that it serves as a unique filter for studying flavor- and spin-dependent contributions to the nucleon structure. The form factor associated with this process is the axial form factor [^1] Also called the nucleon axial-vector form factor; some literature also uses “nucleon transition axial form factor”, since technically it describes a transition from $n\rightarrow p$ or vice versa., $F_{A}(t=-Q^{2})$ [^2] In elastic scattering, the momentum transfer to the proton and the lepton are equivalent. Therefore, $|t|$ and $Q^{2}$ can be used interchangeably, and we use only $|t|$ for convenience in the following text. LlewellynSmith1972. This form factor encodes the momentum-transfer dependence of the nucleon matrix element of the axial current. The so-called nucleon “axial charge radius” [^3] Throughout, we use the term “radius” in the standard form-factor sense, where the radius is defined via the slope of the form factor at $|t|=0$. In a relativistic composite system this need not coincide with a unique, frame-independent three-dimensional static density Miller:2019DefiningRp. , $r_{A}$, is defined via
$\langle r_{A}^{2}\rangle\equiv-\frac{6}{F_{A}(0)}\frac{\mathrm{d}F_{A}}{\mathrm{d}|t|}\big|_{|t|=0},$ (1)
@@
where $F_{A}(0)$ refers to the value of the form factor at zero momentum transfer MINERvA:2023avz. $r_{A}$ encodes the behavior of axial currents globally in the nucleon, and measurement of $r_{A}$ is an attractive goal for future experiments. The authors of Refs. Chen:2024oxx; Chen:2024ksq highlight $r_{A}$ as the primary component of a proton “spin radius”, defined as
$\langle r_{\mathrm{spin}}^{2}\rangle=r_{A}^{2}+\frac{1}{4M^{2}}(1+\frac{2F_{p}(0)}{F_{A}(0)}),$ (2)
where $F_{p}(0)$ is the induced pseudoscalar form factor at zero momentum transfer.
@@
The value of $F_{A}(0)$ is known to high precision from $\beta$-decay processes. The measured value is $F_{A}(0)=-1.2723\pm 0.0023$, which sets the normalization of the form factor. The axial form factor is also a common target of lattice calculations Alexandrou:2017hac; Alexandrou:2020okk; Capitani:2017qpc; Gupta:2017dwj; RQCD:2019jai; Tsuji:2025quu. For the purpose of extracting the axial radius of the nucleon, the necessary quantity to measure is the slope of the form factor as $t\rightarrow 0$, denoted here as $F^{\prime}_{A}(t\rightarrow 0)$. Under the standard assumption that the form factor follows a dipole form, i.e. $F_{A}(t)=F_{A}(0)(1-\frac{t}{M_{A}^{2}})^{-2}$, the slope of the form factor is controlled by the dipole mass parameter $M_{A}$, which can be extracted from the $|t|$-dependence of the scattering cross section (and target-spin asymmetries, as will be discussed later).
@@
Figure 1: Diagram of the CCE process. In the elastic process, unlike charged-current deep inelastic scattering, the $W^{-}$ does not resolve the partonic structure of the nucleon and the process is sensitive to the axial form factor.
@@
The landscape of existing experimental measurements of $F_{A}(t)$ includes quasi-elastic neutrino scattering on nuclei, pion electroproduction on protons, and a measurement of $\bar{\nu_{\mu}}p\rightarrow\mu^{+}n$ on the organic scintillator of the MINERvA experiment Choi:1993PRL_AxialPionElectroprod; Liesenfeld:1999PLB_AxialFromPeepiN; T2K:2020CC0piOC; AguilarArevalo:2010MiniBooNECCQE; MINERvA:2023avz. Both the neutrino data with nuclear targets and the pion electroproduction data suffer from large epistemic uncertainties and model dependence NuSTEC:2017hzk; Haberzettl:2000sm [^4] A recent follow-up article from the MINERvA collaboration further highlights this point Meyer:2025rzh.. The MINERvA data therefore provide the cleanest existing measurement of the axial form factor, and the value of $r_{A}$ extracted from that data is $0.73\pm 0.17$ fm, which is consistent with the charge radius of the proton within the uncertainties.
@@
The goal of this manuscript is to investigate the feasibility of measuring the CCE process at the EIC. The highly polarized proton beam at the EIC provides the opportunity to gain novel information about the axial form factor via the target-spin asymmetries of the CCE process  [^5] Throughout this paper, “target” in the context of the EIC always refers to the proton beam, as is standard parlance in hadronic physics. Target-spin asymmetries thus refer to differences in the process cross section between different proton polarization directions. This is despite the fact that at the EIC there is nominally no fixed target in the traditional sense (see, however, Ref. CFNS2025FixedTargetEIC).. A polarized proton target is generally infeasible in neutrino scattering experiments due to the large target mass required to acquire adequate statistics for a measurement. Refs. Tomalak:2020zlv; Graczyk:2019xwg calculate the target-spin asymmetries for the crossed process $\nu n\rightarrow l^{-}p$ for both longitudinal and transverse target polarizations. These target-spin asymmetries are calculated to be very large, the longitudinal asymmetry $A_{UL}$ reaches nearly 100% at low $Q^{2}$ and the transverse asymmetry $A_{UT}$ is greater than 50% for $Q^{2}>0.5\mathrm{~GeV}^{2}$. These asymmetries have never been measured, and it seems unlikely that they can be measured on proton targets at any existing or proposed facility, with the possible exception of the EIC. The $|t|$-dependence of these target-spin asymmetries furthermore encodes information about the $|t|$-dependence of $F_{A}$, and therefore they can be used to augment the extraction of $r_{A}$ from scattering cross sections Graczyk:2019xwg. A combined fit of $F_{A}(t)$ to the target-spin asymmetry and $|t|$-dependence of the cross section of the CCE process in $ep$ scattering would thus maximize the information about the form factor that could be obtained at the EIC.
@@ II Measurement of Charged-Current Elastic Scattering at the EIC
Figure 2: Compilation of data for the reaction $\nu_{\mu}n\rightarrow\mu^{-}p$ reproduced from Ref. Formaggio:2012cpf. By crossing symmetry, this cross section should be the same as $e^{-}p\rightarrow\nu_{e}n$, neglecting subleading effects such as lepton masses and higher-order corrections.
@@
Figure 4: Expected event yield in 100 fb-1 of integrated luminosity for the EIC in the 10x275 GeV configuration. The bin width in $|t|$ is 0.4 GeV2.
@@ V Projections
Figure 7: Left: Projected yields as a function of $|t|$ for 500 fb-1. Center: Background-subtracted CCE cross section. Right: Projected target-single-spin asymmetry including uncertainty on the background asymmetry $A_{bg}$. See text for more details.
@@
Assuming a veto system capable of reducing the number of background events to 5000 per 100 fb-1, we make projections for the measurement of the CCE cross section and spin asymmetries in Fig. 7. The shape of the background is assumed to have the same functional form as the signal shown in Fig. 4, but with a different dipole mass of 0.8 GeV to avoid bias from having exactly the same background and signal distribution shapes. An integrated luminosity of 500 fb-1, corresponding to 5 years of running at the design luminosity, is used for these projections. The first two panels of Fig. 7 show the total expected yield including background (left) and expected cross section measurement (center) in 500 fb-1 of integrated luminosity. In the background-subtracted cross section projection, the uncertainty is only statistical, taken from the total observed yield in each $|t|$ bin, $\delta N=\sqrt{N_{\rm obs}}$, such that it includes the Poisson fluctuations from both signal and background events. The true data are generated at $M_{A}=1.0$ GeV, and cross section curves corresponding to $M_{A}$ values of 0.8 GeV and 1.2 GeV are overlaid. A precision on $M_{A}$ of around 20% could be expected from a dipole fit to these datapoints. In the dipole ansatz, $r_{A}$ can be expressed solely in terms of $M_{A}$ as $r_{A}^{2}=12M_{A}^{-2}$, such that a 20% uncertainty on $M_{A}$ would correspond to a 20% uncertainty on $r_{A}$ prior to inclusion of target-spin asymmetry data. This precision would be on par with the MINERvA result if it can be realized experimentally, although an excellent understanding of the background and large integrated luminosity would be required to make it a reality.
The right panel of Fig. 7 show projections for the $|t|$-dependent longitudinal target-spin asymmetry. The longitudinal target-spin asymmetry is defined as
$A_{UL}\equiv\frac{1}{P_{p}}\,\frac{N^{\leftarrow}-N^{\rightarrow}}{N^{\leftarrow}+N^{\rightarrow}},$ (4)
where $N^{\leftarrow(\rightarrow)}$ are the event yields for the two proton-helicity states and $P_{p}$ is the average value of the proton longitudinal polarization. The projections for $A_{UL}$ are shown in the right panel of Fig. 7. A value $P_{p}=0.70\pm 0.01$ is assumed for the proton beam polarization, consistent with the EIC expectations Huang:2025gqx; Rathmann:2025jgp; Atoian:2025dib. The proton beam polarization can be either longitudinal or transverse at the interaction point. The degree of proton polarization is monitored via dedicated hadron polarimeters that make use of the left-right asymmetry of the beam scattering on a hydrogen gas jet or carbon foil target AbdulKhalek:2021gbh; Nunes:2022cmc. Propagating the expected $1\%$ proton polarization uncertainty gives a fractional uncertainty $\delta A_{UL}/A_{UL}\simeq\delta P_{p}/P_{p}\approx 1\%$, significantly smaller than the statistical uncertainty for all $|t|$ bins. Thus the beam polarization uncertainty is negligible compared to the projected statistical uncertainties, and in fact a larger polarization uncertainty would be tolerable for this measurement.
@@
The longitudinal target-spin asymmetry of the background process $e\vec{p}\rightarrow n+X$ is currently unknown, but its magnitude will impact how well the CCE asymmetry can be extracted. To account for the target-spin asymmetry of the background that will need to be subtracted, we assume the background has an asymmetry of $0.05\pm 0.025$ and propagate this uncertainty into the final uncertainty on the measured CCE asymmetry. This source of uncertainty can be reduced significantly using even low-luminosity data from the early runs of the EIC to measure the $e\vec{p}\rightarrow n+X$ asymmetry. The datapoints are scattered around a nominal asymmetry of $A_{UL}=1$ according to their total uncertainties to mimic a real measurement. The results for a transverse-spin asymmetry $A_{UT}$ would be the similar to those shown in Fig. 7, barring a very large $A_{UT}$ for the background process $e\vec{p}\rightarrow n+X$. Both transverse and longitudinal proton spin asymmetries can be utilized to constrain $F_{A}(t)$, provided enough integrated luminosity is available.
The results of Ref. Graczyk:2019xwg (in particular Fig. 4 of Ref. Graczyk:2019xwg) show the strong momentum-transfer dependence of the longitudinal target-spin asymmetry on the value of the dipole mass $M_{A}$. These results should be taken with a grain of salt when projecting to the EIC kinematics since the correspond to an incoming neutrino energy of only 1 GeV. Nevertheless, a combined fit of $M_{A}$ to the $|t|$-dependent spin asymmetries and cross section will provide the best constraint, and is a unique possibility offered by the EIC. Finally, the measurement of these spin asymmetries will be valuable generically to cross-check existing results and theoretical understanding with a novel technique [^6] One is reminded of the proton elastic form factor puzzle JeffersonLabHallA:1999epl, where the inclusion of the polarization transfer technique for measuring the nucleon $G_{E}/G_{M}$ form factor ratio revealed a large discrepancy with the standard Rosenbluth method..
@@ VII Conclusion
The charged-current elastic scattering process provides a novel means to study the axial structure of the nucleon and potentially some presently unexplored corners of BSM physics. The recent MINERvA analyses of Refs. MINERvA:2023avz; Meyer:2025rzh have demonstrated the power of utilizing a single nucleon target to accurately constrain the nucleon axial structure. From this perspective, the EIC is well-positioned to contribute to this program via measurement of the CCE process with a polarized proton beam. With its high luminosity and polarized beams of electrons and protons, the EIC is, in principle, an excellent tool for studying this reaction. The cross section for CCE scattering at the EIC is expected to be around 3.5 fb assuming an equal mix of left- and right-handed electrons in the electron beam. Target-spin asymmetries, which can be measured for the first time at the EIC, are predicted to be on the order of 100%, with their magnitudes and $|t|$-dependencies providing additional sensitivity to the axial form factor.
The presence of a dominating background from inclusive neutron photoproduction makes it highly unlikely this process can be measured at the EIC without dedicated instrumentation. If one is optimistic, a dedicated hermetic veto detector may be able to deliver the necessary background suppression, depending on the final implementation of the far-forward region and the level of machine backgrounds. Such a veto system could be incorporated into the far-forward region of a second EIC detector. Even with such a veto system, it will be challenging for the EIC to deliver a competitive measurement of the axial form-factor due to large statistical uncertainties. Leveraging the full information encoded in the target-spin asymmetries and cross section will provide the best constraint on the nucleon axial charge radius $r_{A}$.